OpenStudy (anonymous):

3/x-4+2/x+4=14/(x+4)(x-4)

OpenStudy (anonymous):

Assuming that it's 3/(x-4)+2/(x+4)=14/(x+4)(x-4): put the LHS over a common denominator: [3(x+4) + 2(x-4)] / [(x-4)(x+4)] =14/(x+4)(x-4): 3x + 12 + 2x - 8 = 14 5x = 10 x = 2